For θ = 0°, 30°, 90°, 135°, 270°, and 321° show that sin-θ =-sin θ and tan (π-8) =-tan θ Print the results of θ, the left-hand side, and the right-hand side of each expression for each value of θ. For example, if θ = 73°, the output line might look like 311: for the ta 73.0000000 TAN: for theta3.0000000 right ght ide ide --。.95630 4789 = -3.270 e s 3 2 e left 蠫de --3·270e52 33

need help writing the Fortran code
Answer #1

Here is the FORTRAN code

program trig
REAL x1,x2,x3,x4,x5,x6
REAL y1,y2,y3,y4,y5,y6
y1=0
y2=30
y3=90
y4=135
y5=270
y6=321
z1=180-0
z2=180-30
z3=180-90
z4=180-135
z5=180-275
z6=180-321
  
x1=(0*3.1415927) /180
x2=(30*3.1415927) /180
x3=(90*3.1415927) /180
x4=(135*3.1415927) /180
x5=(270*3.1415927) /180
x6=(321*3.1415927) /180
  
z1=(z1*3.1415927) /180
z2=(z2*3.1415927) /180
z3=(z3*3.1415927) /180
z4=(z4*3.1415927) /180
z5=(z5*3.1415927) /180
z6=(z6*3.1415927) /180
  
PRINT *, "SIN for theta=",y1," left side=", SIN(-x1)," right side=",-SIN(x1)
PRINT *, "TAN for theta=",y1," left side=", TAN(z1)," right side=",-TAN(x1)
  
PRINT *, "SIN for theta=",y2," left side=", SIN(-x2)," right side=",-SIN(x2)
PRINT *, "TAN for theta=",y2," left side=", TAN(z2)," right side=",-TAN(x2)
  
PRINT *, "SIN for theta=",y3," left side=", SIN(-x3)," right side=",-SIN(x3)
PRINT *, "TAN for theta=",y3," left side=", TAN(z3)," right side=",-TAN(x3)
  
PRINT *, "SIN for theta=",y4," left side=", SIN(-x4)," right side=",-SIN(x4)
PRINT *, "TAN for theta=",y4," left side=", TAN(z4)," right side=",-TAN(x4)
  
PRINT *, "SIN for theta=",y5," left side=", SIN(-x5)," right side=",-SIN(x5)
PRINT *, "TAN for theta=",y5," left side=", TAN(z5)," right side=",-TAN(x5)
  
PRINT *, "SIN for theta=",y6," left side=", SIN(-x6)," right side=",-SIN(x6)
PRINT *, "TAN for theta=",y6," left side=", TAN(z6)," right side=",-TAN(x6)
end program trig

As we are dealing with trignomettric funtions there will be slight errors in value but it is very negligible as the left side and right side values will be too close